Product Cost Concept of Product Pricing

Mademoiselle Company produces women's handbags. The cost of producing 1,180 handbags is as follows:

Direct materials $14,000
Direct labor 8,200
Factory overhead 6,500
Total manufacturing cost $28,700

The selling and administrative expenses are $26,300. The management desires a profit equal to 15% of invested assets of $503,000.

If required, round your answers to nearest whole number.

a. Determine the amount of desired profit from the production and sale of 1,180 handbags.
$

b. Determine the product cost per unit for the production of 1,180 handbags.
$per unit

c. Determine the product cost markup percentage for handbags.
%

d. Determine the selling price of handbags. Round your answers to nearest whole value.

Cost $per unit
Markup $per unit
Selling price $per unit
Answer #1
A. Determination the amount of desired profit from the production and sale of 1,180 handbags
Invested assets 503000
Desired Profit percentage 15%
Desired Profit (503000*15%) 75450
B. Determination the product cost per unit for the production of 1,180 handbags.
Total manufacturing cost 28700
Number of handbags 1180
Product cost per unit (28700/1180) 24
C. Determination the product cost markup percentage for handbags.
Profit = Sales - Manufacturing costs - selling and administrative expenses
75450 = x*1180 - 28700 - 26300
75450+26300+28700 = X*1180
130450 = X*1180
X=130450/1180
X = 110.551 or 111
Selling Price = Product cost *(1+Markup)
110.551 = 24.32203*(1+markup)
110.551/24.32203 = (1 + Markup)
4.545296 = 1 + Markup
Markup = 4.545296 -1
Markup = 3.545296 or 355%
D. Determination of the selling price of handbags
Cost 24
Markup 355%
Selling Price (24.32203*(1+355%)) 111